# Replacing Paths with Connection-Biased Attention for Knowledge Graph Completion

### v1

@@ Introduction
Knowledge graphs (KGs) store facts expressed in the forms of relationships between entities. Each fact is represented as a triple of (h,r,t) or (head, relation, tail). Here, the head and tail represent entities such as people, places, and institutions, and the relation represents the relation between the two entities. These facts are modeled as a directed graph with labeled edges, where each entity is a vertex in the graph, the relations are the edge labels, and edges are directed from the head to the tail entities.
@@
Figure 1: Example of training graph and test queries for KG completion in transductive and inductive settings. The training graph consists of facts regarding movie adaptations of books. The inductive query includes unseen entities and the test graph provides context about these new entities. Transductive query involves entities seen during training.
@@
Real-world KGs grow continuously as new facts are added, and the set of entities may grow over time, meaning that at any given point, completion models may need to be used on entities that were not seen during training. Inductive KG Completion assumes the relations between entities remain the same in the newly added facts. To illustrate this setting, Figure 1 depicts a training graph that reflects the correlation between the role of a director and screenwriter. The same artist often plays these two roles when the screenplay is adapted from an existing novel. In the transductive KG completion setting, a relation prediction task involves queries where both entities are present in the training graph and the most plausible relation is to be determined. In the inductive setting, the test graph contains entities added to the graph after the model was trained, and the inductive KG completion task to find the best option for a missing entity involves entities unseen during training (marked by purple ink in the figure).
@@
Naturally, in the inductive setting, graph-learning-based approaches rely greatly on the information provided by neighbors of the incomplete triple of interest. This has led to several Graph Convolution-based models that encode the surrounding subgraph to learn about relation interactions. Lately, paths between entities have also been utilized in models (Lin et al. 2022; Li, Wang, and Mao 2023; Pan et al. 2022; Zhu et al. 2021), as explicitly encoding path information has resulted in a great performance boost. Here, path information is usually defined as an ordered sequence of relations between two entities. For example, two paths between Stanley Kubrick and Stephen King, according to Figure 1, are $(screenwriter^{-1},adaptedFrom,author)$ and $(director^{-1},adaptedFrom,author)$. One could also find a path of length 6 between Stanley Kubrick and Frank Darabont.
@@
The use of paths necessitates multiple decisions, such as determining the ideal path length to obtain meaningful information and exclude noise, the number of paths to be extracted, and various design choices for combining this information with the subgraph information. Additionally, path extraction between entities on the fly and representing them in the model adds overheads to the training time and parameter count.
@@
Transformers have replaced RNNs for sequence modeling precisely because appropriate positional encoding allows attention alone to suffice; we hypothesize that when represented as sequences of triples, appropriate positional encodings similarly unleash the full power of attention, obviating the need for explicit and costly modeling of path information. Indeed, by adapting the connection-biased attention from GRAN (Wang et al. 2021)– where it was used to better represent single $n$-ary facts in a knowledge base– to the representation of subgraphs of a KG, we demonstrate that Transformers alone, without specialized submodules for path representation, suffice to perform information fusion for accurate KG completion.
@@
Knowledge graph completion is often based on learning continuous vector embedding of entities and relations. KG completion garnered attention in the early 2010s with TransE (Bordes et al. 2013) where the distance between learned embedding vectors determined the plausibility of a triple. GraIL (Teru, Denis, and Hamilton 2020) introduced the inductive learning task and the standard benchmark datasets for this task and has garnered much attention in the last 5 years as a more practical approach to learning for this task. In this section, we discuss the development of both settings.
@@
Initial work in the transductive setting focused on triple-based models, exploiting entities’ inherent properties. Recent work has focused on gathering information about the neighborhood surrounding an entity. This section overviews the baseline models relevant to our transductive relation prediction experiment.
@@
Another family of models aims to capture semantic similarity by observing pairwise interactions between entities. RESCAL (Nickel, Tresp, and Kriegel 2011) was an early model with this idea. It was followed by a simplified variation DistMult (Yang et al. 2015) and a complex number variation ComplEx (Trouillon et al. 2016). SimplE (Kazemi and Poole 2018) extends Canonical Polyadic (CP) decomposition by removing the independence between learning entity representations.
@@
Inductive KG completion aims to extend the task to unseen entities. Models first relied on GNNs and more recently on Transformers to encode context information by aggregating neighborhood interactions. Many models have leveraged path information between the head and tail entities by incorporating a path encoding module in addition to the subgraph module that aggregates neighborhood interactions.
@@
GraIL selects the common neighbors of the target head and tail entity as the context of a target triple to be scored. The model employs double-radius labeling (distance from head, distance from tail) to denote each entity’s relative position and learns embeddings for relations through attention computation inside an R-GCN module. The final score is a function of the triple in consideration and the encoded subgraph surrounding it. CoMPILE (Mai et al. 2021) extends GraIL by considering directedness in its subgraph encoding module and computing edge (triple) attentions to learn edge embedding. TACT (Chen et al. 2021) expands GraIL by proposing an additional relation correlation module by learning 6 predefined categories of interactions via unique linear transformations for each kind. This categorization of TACT is closely related to our approach to constructing a connection-biased adjacency matrix of 7 categories. However, TACT incorporates this information into its subgraph encoding module whereas we use it to compute connection-biased attention in the transformer layers.
@@
NBFNet (Zhu et al. 2021) offers a more scalable solution to this by generalizing Bellman-Ford algorithm for finding the shortest paths. The model learns entity pair representations as well as path representations utilizing the distributive properties of the generalized operators. This allows parallel scoring of query triples that share the same entity-relation pairs and the models suffer from time complexity as well as memory overhead due to additional support needed to represent paths.
While these models perform better than those without path information, they come with a drawback of real-time path extraction for inference triples. Since the core of inductive link prediction is to conduct reasoning in unseen entities during training, preprocessing of paths is not a realistic choice for inference triples.
@@
DRUM (Sadeghian et al. 2019) and NeuralLP (Yang, Yang, and Cohen 2017) extract probabilistic first-order logic rules to assign weights to paths between entities. RuleN (Meilicke et al. 2018) assigns confidence to rules by randomized process. While these methods are efficient in learning short and simple rules, they suffer from scalability issues while finding frequent patterns in large graphs.
@@
    Test-time inference triples: $\mathcal{F}_{\mathrm{infer}}\subset(\mathcal{E}_{\mathrm{test}}\times\mathcal{R}\times\mathcal{E}_{\mathrm{test}})$. Given an incomplete fact from $\mathcal{F}_{\mathrm{infer}}$, the aim is to complete it using information from $\mathcal{G}_{\mathrm{test}}$ and the model trained on $\mathcal{G}_{\mathrm{train}}$.
The inductive setting is characterized by the fact that $\mathcal{E}_{\mathrm{train}}\cap\mathcal{E}_{\mathrm{test}}=\varnothing$, i.e. the entities seen at test time were unseen at training time.
Our goal is to find a model that computes a score $s$ for a triple $\langle h,r,t\rangle$. We hypothesize that a plausibility score of a triple can be determined using information present in the ego graphs of the head and tail entities. The $k$-hop ego graph $\mathcal{N}_{e}$ of entity $e$ consists of the triples in its $k$-hop enclosing subgraph. Thus we model the score with
@@
In the transductive setting, there is no separate test graph and the test-time inference triples satisfy $F_{\mathrm{infer}}\subset(\mathcal{E}_{\mathrm{train}}\times\mathcal{R}\times\mathcal{E}_{\mathrm{train}})$. That is, the entities seen at test time are all present in the training graph. Our goal in relation prediction is to, given putative head and tail entities, predict the relation between them. There are typically many fewer relations than entities (a few hundred vs tens of thousands), so instead of a scoring function, we explicitly model the likelihood of each relation given the putative head and tail entities:
@@
This section describes the architecture used for $g$ in the CBLiP model. Given $\langle h,r,t\rangle$, we find the neighboring triples of both $h$ and $t$ as the context of that triple. In the inductive setting, each triple in our model is represented using learned entity-role vectors for its head and tail entity along with a relation vector embedding. The resulting vector encodings of the triples in the neighborhood are used as inputs to a connection-biased Transformer, and the output sequence is passed through a linear transformation to obtain the final score. Below, we explain these components in detail and the modifications made for the transductive setting.
Entity-role vectors
@@
Here, $\mathbf{f}^{\star}$ is the embedding of the putative completed triple that is to be scored. This embedding is additionally aggregated (using $\oplus$) with a special target vector embedding to distinguish it from the embeddings for triples from $\mathcal{N}$.
@@
Figure 2: An example of constructing a connection-biased adjacency matrix. The colored squares denote the presence of a particular kind of overlap of entities between triples. Cells are left blank for when the triples share neither head nor tail.
Paths are often used to learn relation interaction patterns in a (sub)graph. We want to avoid the design decisions, time, and memory complexity that come with this addition to a model. We aim to learn implicit knowledge of paths, distance, and shared neighborhoods, instrumental to correctly predicting the final relation, by constructing a connection-biased adjacency matrix.
While GRAN (Wang et al. 2021) constructs a similar matrix for each of its $n$-ary fact’s components, we construct this for members in a subgraph. The connection types in our model depict the overlap of entities between neighboring triples whereas GRAN distinguishes connections between $n$-ary fact components such as entity-value, attribute-value, and so on.
@@
This approach serves three main purposes:
  1. 1.
  2. 2.
    It serves as an implicit method of encoding paths as we just need to know about shared entities, and the order of them in a path. For example, we can have a pair of length-2 paths where the involved entities and relations are the same but their directions are different as follows:
    • $\langle e_{1},r_{1},e_{2}\rangle$ and $\langle e_{2},r_{2},e_{3}\rangle$
    • $\langle e_{1},r_{1},e_{2}\rangle$ and $\langle e_{3},r_{2},e_{2}\rangle$
  3. 3.
This implicit knowledge of paths, distance, and shared neighborhood is instrumental to correctly predicting the final relation. We show an example of finding four kinds of connection in figure 2.
Connection-Biased Attention in Transformer Encoder
We denote the output sequence as:
$\mathcal{S}_{out}=[\mathbf{y}^{\star},\mathbf{y}^{1}_{\mathcal{N}},\dots\mathbf{y}^{m}_{\mathcal{N}}]$ (6)
The architecture for connection-biased adjacency is depicted in figure 3 and the overall architecture of the proposed model is depicted in figure 3.
Transductive Training
For using our model in the transductive training, we replace $\text{ROLE}(e)$ with entity-specific learnable vector embeddings. Note that, in this setting, all test entities are seen during training, and a relative representation is not needed. Another change in the transductive version of the model is the distinction between neighbors of head and tail. We choose $m$ neighbors from each entity and construct a specific interpretation of the input token sequence in Equation 3.
@@
Figure 3: Connection-biased attention computation. The left diagram shows a brief overview of the input and output sequence. The diagram on the right elaborates on the proposed enhanced encoder layer with connection bias.
@@
In this section, we evaluate our model on the entity prediction task in the inductive setting and present the performances in three KG datasets (12 versions). Additionally, we present relation prediction results in the transductive setting. All experiments were conducted on Quadro RTX 6000 (with NVLink) GPU with 32 GiB memory.
@@
NeuralLP (Yang, Yang, and Cohen 2017), DRUM (Sadeghian et al. 2019), and RuleN (Meilicke et al. 2018) are rule-extraction-based methods. GraIL (Teru, Denis, and Hamilton 2020), CoMPILE (Mai et al. 2021) and TACT (Chen et al. 2021) are graph-based models that use only the subgraph information surrounding a target entity pair and exclude explicit path information. ConGLR (Lin et al. 2022), Report (Liu et al. 2023), LogCo (Pan et al. 2022), NBFNet (Zhu et al. 2021) are graph-based models that utilize path information between target head and tail entities. Report uses vanilla Transformers to encode context and path and is the most similar to our model in terms of choice of encoding architecture.
@@
DistMult 0.847 0.787 0.891 0.875 0.806 0.936 0.634 0.524 0.720
Models w/o path RotatE 0.799 0.735 0.823 0.970 0.951 0.980 0.729 0.691 0.756
@@
(Teru, Denis, and Hamilton 2020) extracted 12 inductive datasets from three popular KG benchmark datasets- Wordnet, Freebase, and Nell. We present the dataset statistics in supplementary material.
@@
For inductive entity prediction, we corrupt a triple by replacing its head/tail with a randomly chosen entity during training. For inference, we want to see how the model ranks variations of corrupted triples which is done by scoring $\langle h,r,?\rangle$ or ?$,r,t\rangle$ against the full vocabulary and finding the rank of the true triple. However, the subgraph extraction process is necessary yet time-consuming in inductive graph-based learning (Teru, Denis, and Hamilton 2020). We, therefore, follow the existing literature and use 50 randomly chosen entities from $\mathcal{E}_{\mathrm{test}}$ to corrupt test triples. The true triple and the corrupted triples are scored by the model and the rank of the true triple is recorded. The ranks of all true test triples contribute to finding Hits@$n$ (ratio of correct hits in the top $n$ sorted predictions) which is a commonly used metric in rank-based experimental studies. We report Hits@10 and mean reciprocal rank (MRR) and evaluate the performance of our model.
Hyperparameter Selection
@@
The results of the entity prediction task according to Hits@10 are presented in Table 1. Our model CBLiP achieves state-of-the-art performance in 7 dataset splits and the performance is consistently better than the rule-extraction and no-path baselines. The model struggles in the Freebase dataset, where a path-based model best utilizes the extremely rich data density.
The MRR results of this experiment are presented in Table 2, where we notice a similar trend of dominating performance in Wordnet and not in Freebase.
@@
We remove the target relation $r$ from all triples and generate an input sequence $S_{in}$. The output probabilities of each relation type are sorted and the position of true relation $r$ is retrieved. This serves as the rank of the relation. We report Hits@1, Hits@3, and MRR for this task in Table 3. We demonstrate the results in 3 main datasets- namely, WN18RR, FB15K-237, and NELL995. The dataset statistics, hyperparameter configuration, and additional experiments on more datasets can be found in our supplementary material.

### v3

# Sharmishtha Dutta1, Alex Gittens1, Mohammed J. Zaki1, Charu C. Aggarwal2

@@ Introduction
Knowledge graphs (KGs) store facts expressed in the forms of relationships between entities. Each fact is represented as a triple (h,r,t) or (head, relation, tail). Here, the head and tail represent entities such as people, places, and institutions, and the relation represents the relation between the two entities. These facts are modeled as a directed graph with labeled edges, where each entity is a vertex in the graph, the relations are the edge labels, and edges are directed from the head to the tail entities.
@@
Figure 1: Example of training graph and test queries for KG completion in transductive and inductive settings.
@@
Real-world KGs grow continuously as new facts are added, and the set of entities may grow over time, meaning that at any given point, completion models may need to be used on entities that were not seen during training. Inductive KG Completion assumes the relations between entities remain the same in the newly added facts. To illustrate this setting, Figure 1 depicts a training graph that reflects the correlation between the role of a director and screenwriter. The same artist often plays these two roles when the screenplay is adapted from an existing novel. In the transductive KG completion setting, a relation prediction task involves queries where both entities are present in the training graph and the most plausible relation is to be determined. In the inductive setting, the test graph contains entities added to the graph after the model was trained, and the inductive KG completion task to find the best option for a missing entity involves entities unseen during training (marked by green ink in the figure).
@@
Naturally, in the inductive setting, graph-learning-based approaches rely greatly on the information provided by neighbors of the incomplete triple of interest. This has led to several Graph Convolution Network (GCN) based models that encode the surrounding subgraph to learn about relation interactions. Paths between entities have also been utilized in several models (Lin et al. 2022; Li, Wang, and Mao 2023; Pan et al. 2022; Zhu et al. 2021), as explicitly encoding path information has resulted in a great performance boost. Here, path information is usually defined as an ordered sequence of relations between two entities. For example, two paths between Frank D. and Stephen King, according to Figure 1, are $(writer^{-1},adaptedFrom,author)$ and $(director^{-1},adaptedFrom,author)$.
@@
The use of paths necessitates multiple decisions, such as determining the ideal path length to obtain meaningful information and exclude noise, the number of paths to be extracted, and various design choices for combining this information with the subgraph information. Additionally, path extraction between entities on the fly and representing them in the model adds overhead to the training time and parameter count.
@@
Transformers have replaced Recurrent Neural Networks for sequence modeling precisely because appropriate positional encoding allows attention alone to suffice; we hypothesize that when represented as sequences of triples, appropriate positional encodings similarly unleash the full power of attention, obviating the need for explicit and costly modeling of path information. Indeed, by adapting the connection-biased attention from GRAN (Wang et al. 2021)– where it was used to better represent single $n$-ary facts in a knowledge base– to the representation of subgraphs of a KG, we demonstrate that Transformers alone, without specialized submodules for path representation, suffice to perform accurate KG completion.
@@
Knowledge graph completion is often based on learning continuous vector embedding of entities and relations. KG completion garnered attention in the early 2010s with TransE (Bordes et al. 2013) where the distance between learned embedding vectors determined the plausibility of a triple. GraIL (Teru, Denis, and Hamilton 2020) introduced the inductive learning task and the standard benchmark datasets for this task and has garnered much attention as a more practical approach. In this section, we discuss the development of both settings.
@@
Initial work in the transductive setting focused on triple-based models, exploiting entities’ inherent properties. Recent work has focused on gathering information about the neighborhood surrounding an entity.
@@
Another family of models aims to capture semantic similarity by observing pairwise interactions between entities. RESCAL (Nickel, Tresp, and Kriegel 2011) was an early model with this idea. It was followed by a simplified variation DistMult (Yang et al. 2015) and a complex number variation ComplEx (Trouillon et al. 2016). SimplE (Kazemi and Poole 2018) extends Canonical Polyadic (CP) decomposition by removing the independence between learned entity representations.
@@
Inductive KG completion aims to extend the task to unseen entities. Models first relied on Graph Neural Networks (GNNs) and more recently on Transformers to encode context information by aggregating neighborhood interactions. Many models have leveraged path information between the head and tail entities by incorporating a path encoding module in addition to the subgraph module that aggregates neighborhood interactions.
@@
GraIL (Teru, Denis, and Hamilton 2020) selects the common neighbors of the target head and tail entities as the context of a target triple to be scored. The model employs double-radius labeling (distance from head, distance from tail) to denote each entity’s relative position and learns embeddings for relations through attention computation inside an Relational GCN module. The final score is a function of the triple in consideration and the encoded subgraph surrounding it. CoMPILE (Mai et al. 2021) extends GraIL by considering directedness in its subgraph encoding module and computing edge (triple) attentions to learn edge embedding. TACT (Chen et al. 2021) expands GraIL by proposing an additional relation correlation module by learning 6 predefined categories of interactions via unique linear transformations for each kind. This categorization of TACT is closely related to our approach to constructing a connection-biased adjacency matrix of 7 categories. However, TACT incorporates this information into its subgraph encoding module whereas we use it to compute connection-biased attention in the transformer layers.
@@
NBFNet (Zhu et al. 2021) offers a more scalable solution to this by generalizing Bellman-Ford algorithm for finding the shortest paths. The model learns entity pair representations as well as path representations utilizing the distributive properties of the generalized operators. This allows parallel scoring of query triples that share the same entity-relation pairs and the models suffer from time complexity as well as memory overhead due to additional support needed to represent paths. While these models perform better than those without path information, they come with a drawback of real-time path extraction for inference triples. Since the core of inductive link prediction is to conduct reasoning in unseen entities during training, preprocessing of paths is not a realistic choice for inference triples.
@@
RPJE (Niu et al. 2020) combines rules and paths by injecting length-2 rules into KG embeddings. DRUM (Sadeghian et al. 2019) and NeuralLP (Yang, Yang, and Cohen 2017) extract probabilistic first-order logic rules to assign weights to paths between entities. RuleN (Meilicke et al. 2018) assigns confidence to rules by randomized process. While these methods are efficient in learning short and simple rules, they suffer from scalability issues while finding frequent patterns in large graphs.
@@
    Test-time inference triples: $\mathcal{F}_{\mathrm{infer}}\subset\mathcal{E}_{\mathrm{test}}\times\mathcal{R}\times\mathcal{E}_{\mathrm{test}}$. Given an incomplete fact from $\mathcal{F}_{\mathrm{infer}}$, the aim is to complete it using information from $\mathcal{G}_{\mathrm{test}}$ and the model trained on $\mathcal{G}_{\mathrm{train}}$.
The inductive setting is characterized by the fact that $\mathcal{E}_{\mathrm{train}}\cap\mathcal{E}_{\mathrm{test}}=\varnothing$, i.e., the entities seen at test time were unseen at training time.
Our goal is to find a model that computes a score $s$ for a triple $\langle h,r,t\rangle$. We hypothesize that a plausibility score of a triple can be determined using information present in the ego graphs of the head and tail entities. The $k$-hop ego graph $\mathcal{N}_{e}$ of entity $e$ consists of the triples in its $k$-hop enclosing subgraph. Thus, we model the score with
@@
In the transductive setting, there is no separate test graph and the test-time inference triples satisfy $F_{\mathrm{infer}}\subset(\mathcal{E}_{\mathrm{train}}\times\mathcal{R}\times\mathcal{E}_{\mathrm{train}})$. That is, the entities seen at test time are all present in the training graph. Our goal in relation prediction is to, given putative head and tail entities, predict the relation between them. There are typically many fewer relations than entities (a few hundred vs. tens of thousands), so instead of a scoring function, we explicitly model the likelihood of each relation given the putative head and tail entities:
@@
This section describes the architecture used for $g$ in the CBLiP model. Given $\langle h,r,t\rangle$, we find the neighboring triples of both $h$ and $t$ as the context of that triple. In the inductive setting, each triple in our model is represented using learned entity role vectors for its head and tail entity along with a relation vector embedding. The resulting vector encodings of the triples in the neighborhood are used as inputs to a connection-biased Transformer, and the output sequence is passed through a linear transformation to obtain the final score. Below, we explain these components in detail and the modifications made for the transductive setting.
Entity Role Vectors
@@
Here, $\mathbf{f}^{\star}$ is the embedding of the putative completed triple to be scored. This embedding is additionally aggregated (using $\oplus$) with a special target vector embedding to distinguish it from the embeddings for triples from $\mathcal{N}$.
@@
Paths are often used to learn relation interaction patterns in a (sub)graph. We want to avoid the design decisions, time, and memory complexity that come with this addition to a model. By constructing a connection-biased adjacency matrix, we aim to learn implicit knowledge of paths, distance, and shared neighborhoods, which is instrumental in correctly predicting the final relation.
While GRAN (Wang et al. 2021) constructs a similar matrix for each of its $n$-ary fact’s components, we build this for members in a subgraph. The connection types in our model depict the overlap of entities between neighboring triples whereas GRAN distinguishes connections between $n$-ary fact components such as entity-value, attribute-value, and so on.
@@
Figure 2: An example of constructing a connection-biased adjacency matrix. The non-empty cells denote the presence of a particular kind of overlap of entities between triples.
This approach serves three main purposes:
  1. 1.
  2. 2.
    It serves as an implicit method of encoding paths as we just need to know about shared entities, and their order in a path. For example, we can have a pair of length-2 paths where the involved entities and relations are the same but their directions are different as follows:
    • $\langle e_{1},r_{1},e_{2}\rangle$ and $\langle e_{2},r_{2},e_{3}\rangle$
    • $\langle e_{1},r_{1},e_{2}\rangle$ and $\langle e_{3},r_{2},e_{2}\rangle$
  3. 3.
This implicit knowledge of paths, distance, and shared neighborhood is instrumental to correctly predicting the final relation. Figure 2 shows an example of finding four kinds of connections.
Connection-Biased Attention in Transformer Encoder
We denote the output sequence as:
$\mathcal{S}_{out}=[\mathbf{y}^{\star},\mathbf{y}^{1}_{\mathcal{N}},\dots,\mathbf{y}^{m}_{\mathcal{N}}]$ (6)
The architecture for connection-biased adjacency and the overall architecture of the proposed model is depicted in Figure 3.
Transductive Training
For using our model in the transductive training, we replace $\text{ROLE}(e)$ with entity-specific learnable vector embeddings. Note that, in this setting, all test entities are seen during training, and a relative representation is not needed. Another change in the transductive version of the model is the distinction between neighbors of head and tail. We choose $m$ neighbors from each entity and construct a specific interpretation of the input token sequence in Eq. (3).
@@
Figure 3: Connection-biased attention computation. The left diagram shows an overview of the input and output sequence, and the right one elaborates the enhanced encoder layer with connection bias. Here, MatMul* and MatMul** refers to the modified matrix multiplication with key bias and value bias, respectively.
@@
In this section, we evaluate our model on the entity prediction task in the inductive setting and present the performances in three KG datasets (12 versions). Additionally, we present relation prediction results in the transductive setting. All experiments were conducted on Quadro RTX 6000 (with NVLink) GPU with 32 GB memory.
@@
NeuralLP (Yang, Yang, and Cohen 2017), DRUM (Sadeghian et al. 2019), and RuleN (Meilicke et al. 2018) are rule-extraction-based methods. GraIL (Teru, Denis, and Hamilton 2020), CoMPILE (Mai et al. 2021) and TACT (Chen et al. 2021) are graph-based models that use only the subgraph information surrounding a target entity pair and exclude explicit path information. ConGLR (Lin et al. 2022), Report (Liu et al. 2023), LogCo (Pan et al. 2022), NBFNet (Zhu et al. 2021) are graph-based models that utilize path information between target head and tail entities. Report (Li, Wang, and Mao 2023) uses vanilla Transformers to encode context and path and is the most similar to our model in terms of choice of encoding architecture.
@@
Models w/o path DistMult 0.847 0.787 0.891 0.875 0.806 0.936 0.634 0.524 0.720
RotatE 0.799 0.735 0.823 0.970 0.951 0.980 0.729 0.691 0.756
@@
(Teru, Denis, and Hamilton 2020) extracted 12 inductive datasets from three popular KG benchmark datasets – Wordnet, Freebase, and Nell. We present the dataset statistics in supplementary material (Dutta et al. 2024).
@@
For inductive entity prediction, we corrupt a triple by replacing its head/tail with a randomly chosen entity during training. For inference, we want to see how the model ranks variations of corrupted triples by scoring $\langle h,r,?\rangle$ or ?$,r,t\rangle$ and finding the rank of the true triple. We follow the existing literature and use 50 randomly chosen entities from $\mathcal{E}_{\mathrm{test}}$ to corrupt test triples. The model scores the true triples and the corrupted triples, and the rank of the true triple is recorded. The ranks of all true test triples contribute to finding Hits@$n$ (ratio of correct hits in the top $n$ sorted predictions), which is a commonly used metric in rank-based experimental studies. We report Hits@10 and mean reciprocal rank (MRR) and evaluate the performance of our model. The hyperparameter selection is described in the supplementary material (Dutta et al. 2024).
@@
The results of the entity prediction task according to Hits@10 are presented in Table 1. Our model CBLiP achieves state-of-the-art performance in 7 dataset splits and the performance is consistently better than the rule-extraction and no-path baselines. The model struggles in the Freebase dataset, where a path-based model best utilizes the extremely rich data density. The MRR results of this experiment are presented in Table 2, where we notice a similar trend of dominating performance in Wordnet and not in Freebase.
@@
We remove the target relation $r$ from all triples and generate an input sequence $S_{in}$. The output probabilities of each relation type are sorted and the position of true relation $r$ is retrieved. This serves as the rank of the relation. We report Hits@1, Hits@3, and MRR for this task in Table 3. We demonstrate the results on three primary datasets: WN18RR, FB15K-237, and NELL995. Our supplementary material (Dutta et al. 2024) contains the dataset statistics, hyperparameter configuration, and additional experiments on other datasets.